# LittleBit: Ultra Low-Bit Quantization via Latent Factorization

### v1

@@ 4.2 Main Results
Table 1 shows LittleBit’s perplexity (PPL) results against baselines. LittleBit consistently performs robustly, especially in the sub-1-bit regime. Compared to STBLLM [16], LittleBit achieves markedly superior PPL at 0.8, 0.7, and 0.55 BPW. For instance, on Llama2-7B, LittleBit at 0.55 BPW achieves a PPL of 10.47, a substantial improvement over STBLLM’s 30.67. Similar significant gains are observed across other models and BPWs; for Llama2-13B at 0.8 BPW, LittleBit records a PPL of 8.52 compared to STBLLM’s 11.90.
@@
LittleBit’s advantages are consistently observed across a diverse set of model architectures and sizes. This includes the Llama3 family, which is often considered challenging for quantization, and other contemporary models like Phi-4. For example, Llama3-8B quantized by LittleBit to 0.55 BPW yields a PPL of 18.47, significantly outperforming STBLLM’s 241.95 (at 0.55 BPW with 4:8 sparsity).The zero-shot evaluation results presented in Table 2 further reinforce these perplexity-based findings. For instance, Llama2-7B compressed by LittleBit to 0.55 BPW achieves a mean accuracy of 47.26%, and even at an extreme 0.3 BPW, it maintains 45.20% mean accuracy. These results are notably competitive with, and in several cases surpass, those of STBLLM at similar or even higher effective bits (e.g., STBLLM Llama2-7B at 0.55 BPW scores 44.29%). This suggests a superior preservation of the models’ intrinsic reasoning capabilities even under extreme compression, highlighting LittleBit’s efficacy in creating highly compact yet powerful LLMs.
@@ 5 Analysis
Designed for resource-constrained environments such as on-device deployment, LittleBit achieves substantial reductions in model memory footprint. As summarized in Table 3, applying LittleBit to quantize a Llama2-7B model (originally 13.49 GB in FP16) to an effective bit of 0.3 BPW reduces its required storage to 0.79 GB. This corresponds to a compression factor exceeding 17$\times$. At the extreme setting of 0.1 BPW, the footprint is further reduced to 0.63 GB (over 21$\times$ compression). Comparable reductions are observed for larger models such as Llama2 70B, where the memory footprint decreases from 138.04 GB to under 2 GB at 0.1 BPW, achieving nearly 70$\times$ compression. These substantial memory reductions significantly expand the range of devices capable of hosting large-scale language models, thereby enabling complex language tasks on hardware with limited VRAM or storage capacity.
Table 3: Memory footprint comparison (GB) for Llama2 [49] models under different quantization methods. LittleBit is evaluated at various BPWs, achieved by adjusting the latent rank $r$. Compression factors relative to FP16 are shown in parentheses.
@@
Effective Bit 16 1.1 1 0.8 0.55 0.3 0.1
Llama2-7B 13.49 GB 1.60 GB (8.43$\times$) 1.36 GB (9.92$\times$) 1.19 GB (11.34$\times$) 0.98 GB (13.77$\times$) 0.79 GB (17.08$\times$) 0.63 GB (21.41$\times$)
Llama2-13B 26.06 GB 2.80 GB (9.31$\times$) 2.28 GB (11.43$\times$) 1.95 GB (13.36$\times$) 1.51 GB (17.26$\times$) 1.15 GB (22.66$\times$) 0.84 GB (31.02$\times$)
Llama2-70B 138.04 GB 15.40 GB (8.96$\times$) 9.72 GB (14.20$\times$) 7.97 GB (17.31$\times$) 5.83 GB (23.68$\times$) 3.70 GB (37.31$\times$) 1.98 GB (69.72$\times$)
@@
Figure 7: Latency (ms) on NVIDIA A100 (1 batch) for Llama2-70B MLP (8,192$\times$28,672). Compares native PyTorch FP16 GEMM with OneBit and LittleBit (LB) custom kernels at 1.0, 0.55, 0.3, and 0.1 BPW.
@@
Beyond memory savings, inference speed is crucial. LittleBit’s factorized architecture (Eq. 5), unlike memory-focused or hardware-dependent methods, targets computational speedup via low-rank binary operations. While introducing an extra latent scaling step, we hypothesized that the reduced computational cost from binary operations at low ranks r would decrease overall latency. To assess this, we benchmarked a custom 1-bit GEMV CUDA kernel (details in Section A.8) focusing on the primary path computation. As shown in Fig. 7 for a Llama2-70B MLP layer, LittleBit’s latency is comparable to FP16 near 1.0 BPW but achieves substantial speedups (up to 5.5$\times$ relative to FP16) in the sub-0.5 BPW regime. These preliminary results suggest LittleBit’s potential for efficient deployment with significant latency reduction, especially at extreme compression levels.
@@ 6 Conclusion
In this work, we proposed LittleBit, a novel method advancing LLM compression to the extreme sub-0.5 BPW regime, with viable performance at 0.1 BPW. Gains were achieved via SVD-inspired latent matrix factorization, binarization of these factors, and a multi-scale (row, column, latent) compensation mechanism. We also introduced Dual-SVID initialization for stable quantization-aware training (QAT) and integrated Residual Compensation for error mitigation. Empirical results show LittleBit significantly outperforms prior sub-1-bit techniques and maintains robust fidelity across LLMs up to 32B parameters. LittleBit thus establishes a new state-of-the-art in the size–performance trade-off, significantly expanding LLM deployment in resource-constrained environments. Future research will explore hardware co-design and further extensions towards the ultra-low bit frontier.
@@ A.1 Mathematical Analysis and Proofs
As stated in Section 3.1 (Proposition 1), for input $\mathbf{X}\in\mathbb{R}^{\mathrm{seq}\times d_{\mathrm{in}}}$, the primary quantized weight matrix is $\widehat{\mathbf{W}}_{\mathrm{pri}}=\mathrm{diag}(\mathbf{h})\mathbf{U}_{\mathrm{sign}}\mathrm{diag}(\bm{\ell})\mathbf{V}_{\mathrm{sign}}^{\top}\mathrm{diag}(\mathbf{g})$. The forward pass $\mathbf{Y}=\mathbf{X}\widehat{\mathbf{W}}_{\mathrm{pri}}^{\top}$ can be computed as shown in Eq. 12.
@@
Terms are defined in Section 3.1.
@@
This yields Eq. 12, where $\odot$ implies element-wise multiplication with broadcasting. ∎
Lemma 1 (Quantization Error vs. Factor Rank).
@@
Lemma 2 (Quantization Bias vs. SVD Component Structure).
@@
Theorem 1 (Potential Advantage of Two-Stage Quantization of SVD Components).
Using the setup of Lemma 2, let $\widehat{\mathbf{W}}_{r}^{\mathrm{single}}=\mathcal{Q}(\widetilde{\mathbf{W}}_{r})$ and $\widehat{\mathbf{W}}^{\mathrm{two-stage}}=\mathcal{Q}(\widetilde{\mathbf{W}}_{r_{1}})+\mathcal{Q}(\widetilde{\mathbf{R}}_{r_{2}})$. The two-stage error $\mathcal{E}_{\mathrm{two-stage}}=\left\|\mathbf{W}-\widehat{\mathbf{W}}^{\mathrm{two-stage}}\right\|_{F}$ can be less than the single-stage error $\mathcal{E}_{\mathrm{single-stage}}=\left\|\mathbf{W}-\widehat{\mathbf{W}}_{r}^{\mathrm{single}}\right\|_{F}$. This occurs if the quantization error of the sum is greater than the sum of (vectorial) quantization errors of parts:
@@
Let $\mathbf{E}_{\mathrm{trunc}}=\mathbf{W}-\widetilde{\mathbf{W}}_{r}$. Errors are $\Delta_{1}=\mathcal{Q}(\widetilde{\mathbf{W}}_{r_{1}})-\widetilde{\mathbf{W}}_{r_{1}}$, $\Delta_{2}=\mathcal{Q}(\widetilde{\mathbf{R}}_{r_{2}})-\widetilde{\mathbf{R}}_{r_{2}}$, $\Delta_{\mathrm{sum}}=\mathcal{Q}(\widetilde{\mathbf{W}}_{r})-\widetilde{\mathbf{W}}_{r}$. Then $\mathcal{E}_{\mathrm{two-stage}}=\|\mathbf{E}_{\mathrm{trunc}}-(\Delta_{1}+\Delta_{2})\|_{F}$ and $\mathcal{E}_{\mathrm{single-stage}}=\|\mathbf{E}_{\mathrm{trunc}}-\Delta_{\mathrm{sum}}\|_{F}$. The condition Eq. 15 makes $\mathcal{E}_{\mathrm{two-stage}}<\mathcal{E}_{\mathrm{single-stage}}$ likely. Non-linear $\mathcal{Q}$ (e.g., with adaptive scaling) can satisfy this by better handling the distinct characteristics of $\widetilde{\mathbf{W}}_{r_{1}}$ and $\widetilde{\mathbf{R}}_{r_{2}}$. This motivates LittleBit’s separate residual handling, though its residual is learned differently. ∎
@@ A.2.1 Residual Compensation
As shown in  Fig. 8, residual compensation preserves a high-fidelity resemblance to the full-precision baseline across all inspected layers. In the no-residual setting, the output feature distributions exhibit noticeable deformation relative to the full-precision baseline. In contrast, applying residual compensation consistently restores structural similarity to the original distribution, preserving both directional patterns and magnitude diversity.
Table 5 presents the PPL results for the OPT-1.3B model, comparing configurations with and without residual compensation. As indicated in the table, at the extremely low bit of 0.1 BPW, the version with residual compensation (PPL 60.011) performed worse than the version without it (PPL 48.512). This suggests that for relatively small models (1.3B), when pushing effective bits to such extreme lows, the additional complexity introduced by the residual compensation mechanism might outweigh its benefits and instead hinder performance. However, for the same OPT-1.3B model at other effective bits evaluated (ranging from 0.3 BPW to 1.0 BPW), the inclusion of residual compensation consistently resulted in better (lower) PPL. Moreover, experiments conducted with other, larger models generally shown that residual compensation tended to provide superior performance. In light of these overall advantages observed across various settings and particularly for larger models, residual compensation was adopted as a standard component in all experiments presented in this study.
@@ A.2.2 SmoothSign versus Straight-Through Estimator
For backpropagating through the non-differentiable $\mathrm{sign}(x)$ function, we compared the Straight-Through Estimator (STE) [31] with our proposed SmoothSign technique. SmoothSign employs the $\mathrm{sign}(x)$ function for the forward pass. For the backward pass, it utilizes a smooth proxy gradient, specifically the derivative of $\tanh(kx)$ where $k=100$ (illustrated in Fig. 9).
Table 6 (OPT-1.3B results) shows SmoothSign yielding better PPL as effective bits decrease, with a notable advantage at 0.1 BPW. Due to its superior stability and performance in the ultra-low bit regime, SmoothSign was adopted for all QAT experiments (consistent with Section 4.1).
@@ A.2.3 Grouped Query Attention
To investigate, we performed an ablation on Llama3-8B ($\sim$0.1 BPW target), increasing the latent rank $r$ for K/V layers by 2$\times$, 4$\times$, and 8x over the standard calculation (Section A.4), while other layers remained unchanged. Table 7 shows that a 4x factor for K/V rank yielded the best PPL on WikiText-2 and C4, with minimal impact on overall BPW (e.g., $\sim$0.098 BPW to $\sim$0.107 BPW for 4$\times$). Average zero-shot accuracy also remained competitive. An 8x factor showed diminishing returns.
Given the significant PPL improvement with the 4$\times$ K/V rank factor (nearly 1 point on WikiText-2 vs. 1$\times$) at a negligible BPW cost (<10% relative increase in effective bits for transformer layers, <4% of total parameters), this strategy was adopted for GQA/MQA models (Llama3-8B, Phi-4, QwQ-32B) in our main results. For these models, the latent rank $r$ for K and V projection layers was calculated based on a 4$\times$ multiplier compared to the standard calculation for the target BPW, while all other layers used the standard rank calculation. This approach helps mitigate potential information bottlenecks in GQA/MQA layers under extreme quantization without significantly altering the target compression level.
@@ A.3 Perplexity on C4 and PTB
To further assess LittleBit’s generalization, we evaluated its performance on additional benchmarks: the diverse C4 web text corpus and the distinct PTB news corpus (Table 8). Focusing on Llama2-7B/13B against STBLLM, the results on these datasets confirm the trends observed on WikiText-2. LittleBit consistently achieves lower perplexity than STBLLM at comparable sub-1-bit regimes. This performance advantage becomes more significant below 0.55 BPW, where STBLLM’s performance degrades considerably, particularly on PTB.
@@ A.4 Average Bits Per Weight Calculation
This section details the calculation of the average effective bits per weight (which we denote as $b$ in the following equations) for a linear layer implemented using the LittleBit architecture, including the parameters from the primary ($\widehat{\mathbf{W}}_{\mathrm{pri}}$) and residual compensation ($\widehat{\mathbf{W}}_{\mathrm{res}}$) pathways, as described in Sections 3.1 and 3.3. The calculation demonstrates how to determine the latent rank $r$ required to achieve a target value for $b$ within the range of approximately 0.1 to 1.0.
@@
Total Bits $=2\times\left(\text{Bits for }\mathbf{U}_{\mathrm{sign}}+\text{Bits for }\mathbf{V}_{\mathrm{sign}}+\text{Bits for }\mathbf{h}+\text{Bits for}\mathbf{g}+\text{Bits for }\bm{\ell}\right)$
@@
To achieve a target value for $b$, we rearrange Eq. 16 to solve for the required latent rank $r$:
@@
  • Example 1: Linear Layer ($4,096$\times$4,096$$) Let $d_{\mathrm{out}}=$4,096$$ and $d_{\mathrm{in}}=$4,096$$. The original number of parameters is $4,096$\times$4,096$=$16,777,216$$. $d_{\mathrm{out}}+d_{\mathrm{in}}=$8,192$$.
    For a target $b\approx 0.55$: Using Eq. 17 with $b=0.55$:
@@
    We choose integer $r=546$. The actual value of $b$ for $r=546$, using Eq. 16, is:
@@
  • Example 2: Linear Layer ($4,096$\times$11,008$$) Let $d_{\mathrm{out}}=$4,096$$ and $d_{\mathrm{in}}=$11,008$$. The original number of parameters is $4,096$\times$11,008$=$45,090,816$$. $d_{\mathrm{out}}+d_{\mathrm{in}}=$15,104$$.
    For a target $b\approx 0.1$: Using Eq. 17 with $b=0.1$:
@@
    We choose integer $r=133$. The actual value of $b$ for $r=133$, using Eq. 16, is:
@@
By following this procedure for each linear layer in the model, we can select appropriate ranks ($r$) to achieve a desired overall value for $b$ (average bits per weight), thereby controlling the trade-off between model compression and performance. The specific ranks used to achieve the average bits per weight (BPW) figures reported in the main paper (e.g., in Tables 1 and 3) are determined using this calculation methodology for $b$ for each layer type within the respective models.

### v2

@@ 4.2 Main Results
Table˜1 shows LittleBit’s perplexity (PPL) results against baselines. LittleBit consistently performs robustly, especially in the sub-1-bit regime. Compared to STBLLM [16], LittleBit achieves markedly superior PPL at 0.8, 0.7, and 0.55 BPW. For instance, on Llama2-7B, LittleBit at 0.55 BPW achieves a PPL of 10.47, a substantial improvement over STBLLM’s 30.67. Similar significant gains are observed across other models and BPWs; for Llama2-13B at 0.8 BPW, LittleBit records a PPL of 8.52 compared to STBLLM’s 11.90.
@@
LittleBit’s advantages are consistently observed across a diverse set of model architectures and sizes. This includes the Llama3 family, which is often considered challenging for quantization, and other contemporary models like Phi-4. For example, Llama3-8B quantized by LittleBit to 0.55 BPW yields a PPL of 18.47, significantly outperforming STBLLM’s 241.95 (at 0.55 BPW with 4:8 sparsity).The zero-shot evaluation results presented in Table˜2 further reinforce these perplexity-based findings. For instance, Llama2-7B compressed by LittleBit to 0.55 BPW achieves a mean accuracy of 47.26%, and even at an extreme 0.3 BPW, it maintains 45.20% mean accuracy. These results are notably competitive with, and in several cases surpass, those of STBLLM at similar or even higher effective bits (e.g., STBLLM Llama2-7B at 0.55 BPW scores 44.29%). This suggests a superior preservation of the models’ intrinsic reasoning capabilities even under extreme compression, highlighting LittleBit’s efficacy in creating highly compact yet powerful LLMs.
@@ 5 Analysis
[…]
Table 3: Memory footprint comparison (GB) for Llama2 [49] models under different quantization methods. LittleBit is evaluated at various BPWs, achieved by adjusting the latent rank $r$. Compression factors relative to FP16 are shown in parentheses.The gap between the effective BPW and the compression factor arises because quantization is applied only to the Transformer blocks, while components like the embedding layer and the lm_head remain in FP16.
@@
BPW 16 1.1 1 0.8 0.55 0.3 0.1
[…]
@@
[…]
@@
[…]
These results highlight LittleBit’s potential for efficient deployment with significant latency reduction. The observed gap between theoretical computational gains and practical latency stems from the fact that inference at small batch sizes is often dominated by memory access rather than raw computation. Our factorized method involves multiple memory-bound steps, and our kernel, while effective, is not as exhaustively optimized as industry-standard libraries. A detailed analysis including theoretical FLOPs, practical latency, and end-to-end decoding throughput is provided in LABEL:appendix:latency.
@@ 6 Conclusion
In this work, we proposed LittleBit, a novel method advancing LLM compression to the extreme sub-0.5 BPW regime, with viable performance at 0.1 BPW. Gains were achieved via SVD-inspired latent matrix factorization, binarization of these factors, and a multi-scale (row, column, latent) compensation mechanism. We also introduced Dual-SVID initialization for stable quantization-aware training (QAT) and integrated Residual Compensation for error mitigation. Empirical results show LittleBit significantly outperforms prior sub-1-bit techniques and maintains robust fidelity across LLMs up to 32B parameters. LittleBit thus establishes a new state-of-the-art in the size–performance trade-off, significantly expanding LLM deployment in resource-constrained environments. Building on this foundation, future work can enhance practical deployment through hardware co-design for edge devices like neural processing units (NPUs). Further gains in model fidelity can be explored by integrating advanced techniques like non-uniform bit allocation, applied either across different model layers or even within a single layer by dynamically distributing the bit budget between the primary and residual paths.
Acknowledgements
We are grateful to Dr. Heonjae Ha, Dr. SangJeong Lee, Minseop Choi, Changdong Kim, and Hyochan Chong for their insightful discussions and generous support in this work.
@@ A.1 Mathematical Analysis and Proofs
As stated in Section˜3.1 (Proposition 1), for input $\mathbf{X}\in\mathbb{R}^{\mathrm{seq}\times d_{\mathrm{in}}}$, the primary quantized weight matrix is $\widehat{\mathbf{W}}_{\mathrm{pri}}=\mathrm{diag}(\mathbf{h})\mathbf{U}_{\mathrm{sign}}\mathrm{diag}(\bm{\ell})\mathbf{V}_{\mathrm{sign}}^{\top}\mathrm{diag}(\mathbf{g})$. The forward pass $\mathbf{Y}=\mathbf{X}\widehat{\mathbf{W}}_{\mathrm{pri}}^{\top}$ can be computed as shown in Eq.˜12.
@@
Terms are defined in Section˜3.1.
@@
This yields Eq.˜12, where $\odot$ implies element-wise multiplication with broadcasting. ∎
Claim 1 (Quantization Error vs. Factor Rank).
@@
Claim 2 (Quantization Bias vs. SVD Component Structure).
@@
Proposition 2 (Potential Advantage of Two‑Stage Quantization of SVD Components).
Using the setup of ˜2, let $\widehat{\mathbf{W}}_{r}^{\mathrm{single}}=\mathcal{Q}(\widetilde{\mathbf{W}}_{r})$ and $\widehat{\mathbf{W}}^{\mathrm{two-stage}}=\mathcal{Q}(\widetilde{\mathbf{W}}_{r_{1}})+\mathcal{Q}(\widetilde{\mathbf{R}}_{r_{2}})$. The two-stage error $\mathcal{E}_{\mathrm{two-stage}}=\left\|\mathbf{W}-\widehat{\mathbf{W}}^{\mathrm{two-stage}}\right\|_{F}$ can be less than the single-stage error $\mathcal{E}_{\mathrm{single-stage}}=\left\|\mathbf{W}-\widehat{\mathbf{W}}_{r}^{\mathrm{single}}\right\|_{F}$. This occurs if the quantization error of the sum is greater than the sum of (vectorial) quantization errors of parts:
@@
Let $\mathbf{E}_{\mathrm{trunc}}=\mathbf{W}-\widetilde{\mathbf{W}}_{r}$. Errors are $\Delta_{1}=\mathcal{Q}(\widetilde{\mathbf{W}}_{r_{1}})-\widetilde{\mathbf{W}}_{r_{1}}$, $\Delta_{2}=\mathcal{Q}(\widetilde{\mathbf{R}}_{r_{2}})-\widetilde{\mathbf{R}}_{r_{2}}$, $\Delta_{\mathrm{sum}}=\mathcal{Q}(\widetilde{\mathbf{W}}_{r})-\widetilde{\mathbf{W}}_{r}$. Then $\mathcal{E}_{\mathrm{two-stage}}=\|\mathbf{E}_{\mathrm{trunc}}-(\Delta_{1}+\Delta_{2})\|_{F}$ and $\mathcal{E}_{\mathrm{single-stage}}=\|\mathbf{E}_{\mathrm{trunc}}-\Delta_{\mathrm{sum}}\|_{F}$. The condition Eq.˜15 makes $\mathcal{E}_{\mathrm{two-stage}}<\mathcal{E}_{\mathrm{single-stage}}$ likely. Non-linear $\mathcal{Q}$ (e.g., with adaptive scaling) can satisfy this by better handling the distinct characteristics of $\widetilde{\mathbf{W}}_{r_{1}}$ and $\widetilde{\mathbf{R}}_{r_{2}}$. This motivates LittleBit’s separate residual handling, though its residual is learned differently. ∎
@@ A.2.1 Residual Compensation
As shown in  Fig.˜7, residual compensation preserves a high‑fidelity resemblance to the full‑precision baseline across all inspected layers. In the no-residual setting, the output feature distributions exhibit noticeable deformation relative to the full-precision baseline. In contrast, applying residual compensation consistently restores structural similarity to the original distribution, preserving both directional patterns and magnitude diversity.
Table˜5 presents the PPL results for the OPT-1.3B model, comparing configurations with and without residual compensation. As indicated in the table, at the extremely low bit of 0.1 BPW, the version with residual compensation (PPL 60.011) performed worse than the version without it (PPL 48.512). This suggests that for relatively small models (1.3B), when pushing effective bits to such extreme lows, the additional complexity introduced by the residual compensation mechanism might outweigh its benefits and instead hinder performance. However, for the same OPT-1.3B model at other effective bits evaluated (ranging from 0.3 BPW to 1.0 BPW), the inclusion of residual compensation consistently resulted in better (lower) PPL. Moreover, experiments conducted with other, larger models generally showed that residual compensation tended to provide superior performance. In light of these overall advantages observed across various settings and particularly for larger models, residual compensation was adopted as a standard component in all experiments presented in this study.
@@ A.2.2 SmoothSign versus Straight-Through Estimator
For backpropagating through the non-differentiable $\mathrm{sign}(x)$ function, we compared the Straight-Through Estimator (STE) [31] with our proposed SmoothSign technique. SmoothSign employs the $\mathrm{sign}(x)$ function for the forward pass. For the backward pass, it utilizes a smooth proxy gradient, specifically the derivative of $\tanh(kx)$ where $k=100$ (illustrated in Fig.˜8).
Table˜6 (OPT-1.3B results) shows SmoothSign yielding better PPL as effective bits decrease, with a notable advantage at 0.1 BPW. Due to its superior stability and performance in the ultra-low bit regime, SmoothSign was adopted for all QAT experiments (consistent with Section˜4.1).
@@ A.2.3 Grouped Query Attention
[…]
@@ A.3 Perplexity on C4 and PTB
To further assess LittleBit’s generalization, we evaluated its performance on additional benchmarks: the diverse C4 web text corpus and the distinct PTB news corpus (Table˜8). Focusing on Llama2-7B/13B against STBLLM, the results on these datasets confirm the trends observed on WikiText-2. LittleBit consistently achieves lower perplexity than STBLLM at comparable sub-1-bit regimes. This performance advantage becomes more significant below 0.55 BPW, where STBLLM’s performance degrades considerably, particularly on PTB.
@@ A.4 Average Bits Per Weight Calculation
This section details the calculation of the average effective bits per weight (which we denote as $b$ in the following equations) for a linear layer implemented using the LittleBit architecture, including the parameters from the primary ($\widehat{\mathbf{W}}_{\mathrm{pri}}$) and residual compensation ($\widehat{\mathbf{W}}_{\mathrm{res}}$) pathways, as described in Sections˜3.1 and 3.3. The calculation demonstrates how to determine the latent rank $r$ required to achieve a target value for $b$ within the range of approximately 0.1 to 1.0.
@@
[…]
@@
To achieve a target value for $b$, we rearrange Eq.˜16 to solve for the required latent rank $r$:
@@
  • Example 1: Linear Layer ($4,096$\times$4,096$$) Let $d_{\mathrm{out}}=$4,096$\$ and $d_{\mathrm{in}}=$4,096$\$. The original number of parameters is $4,096$\times$4,096$=$16,777,216$$. $d_{\mathrm{out}}+d_{\mathrm{in}}=$8,192$\$.
    For a target $b\approx 0.55$: Using Eq.˜17 with $b=0.55$:
@@
    We choose integer $r=546$. The actual value of $b$ for $r=546$, using Eq.˜16, is:
@@
  • Example 2: Linear Layer ($4,096$\times$11,008$$) Let $d_{\mathrm{out}}=$4,096$\$ and $d_{\mathrm{in}}=$11,008$\$. The original number of parameters is $4,096$\times$11,008$=$45,090,816$$. $d_{\mathrm{out}}+d_{\mathrm{in}}=$15,104$\$.
    For a target $b\approx 0.1$: Using Eq.˜17 with $b=0.1$:
@@
    We choose integer $r=133$. The actual value of $b$ for $r=133$, using Eq.˜16, is:
@@
By following this procedure for each linear layer in the model, we can select appropriate ranks ($r$) to achieve a desired overall value for $b$ (average bits per weight), thereby controlling the trade-off between model compression and performance. The specific ranks used to achieve the average bits per weight (BPW) figures reported in the main paper (e.g., in Tables˜1 and 3) are determined using this calculation methodology for $b$ for each layer type within the respective models.